# Adapting Vision-Language Models for Evaluating World Models

### v1

@@ D.1.3 Prompt Format
$S=\;$ $\underbrace{\texttt{<image>}_{1}^{(1)},\ldots,\texttt{<image>}_{p}^{(1)},\;\ldots,\;\texttt{<image>}_{1}^{(n)},\ldots,\texttt{<image>}_{p}^{(n)}}_{S_{\mathcal{I}}\text{: Visual tokens from }n\text{ frames, each represented as }p\text{ patches}}$
@@
D.1.4 Pretraining Data and Filtering
PaliGemma is pretrained on a mixture of large-scale vision-language datasets, including WebLI [23], CC3M-35L [87], VQ2A-CC3M-35L [19], OpenImages [80], and WIT [90]. Data quality and safety are maintained through pornographic content filtering, text safety and toxicity filtering, and privacy-preserving measures.
D.2 Evaluation Metrics
@@
Formally, let $G$ and $R$ denote the sets of $n$-grams in the generated and reference answers, respectively. Precision and recall are defined as:
$P=\frac{|G\cap R|}{|G|},\quad R=\frac{|G\cap R|}{|R|}$ (3)
where $|G\cap R|$ counts overlapping $n$-grams. The ROUGE score is then computed as:
@@ D.3 Results
Optimizer AdamW [67]
@@
VLM backbone PaliGemma 2 (3B) [10]
Hyperparameters. Table 3 summarizes the core training hyperparameters used across all adaptation experiments. We train all models on 8 NVIDIA A100 GPUs with a batch size of 1 per device and accumulate gradients over 4 steps, yielding an effective batch size of 32. Each epoch corresponds to a full pass over the adaptation dataset, and no early stopping is applied. Models were trained for 1–10 epochs depending on task and setting. Optimization is performed using AdamW [67] with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, a base learning rate of $5\times 10^{-5}$, and weight decay of $1\times 10^{-6}$. We use cosine learning rate annealing [68] with a linear warmup over the first 10% of training steps. To stabilize training, we apply gradient clipping with a global norm threshold of 1.0. All models use PaliGemma 2 (3B) [10] as the vision-language backbone unless otherwise noted. We vary the number of input frames between 1 and 8 depending on task, and all images are resized to a fixed resolution of $224\times 224$. Training is conducted in bfloat16 precision using data parallelism. Model selection is based on final validation accuracy.
@@
Appendix E Human Annotation Study
@@ E.1 Study Design
Selected World Models. For our study, we select two autoregressive world models of different scales, both based on the WHAM architecture [56], a publicly available world model. These models are trained to model sequences of visual frames and controller actions, without any textual supervision. Each world model is a decoder-only transformer [81, 98] trained to autoregressively predict discrete tokens representing visual observations and actions. Visual frames are first encoded using a VQGAN [32], while joystick actions are tokenized using a learned discretization scheme based on action bucketization [55]. The model is trained to predict the next token in the sequence, conditioned on prior visual and action tokens. Specifically, we focus on two versions of WHAM with differing model capacities and training environments:
  • WHAM 140M: A 140M-parameter model trained for 100K steps on gameplay from a single environment (Environment A / Skygarden) at $128\times 128$ resolution.
  • WHAM 1.6B: A 1.6B-parameter model trained for 200K steps on gameplay from seven environments (Environments A–G, including Skygarden) at $300\times 180$ resolution.
Rollouts Generation. Rollout generation follows a consistent protocol for both world models: at inference time, the model is conditioned on 1 second of ground-truth gameplay (visual and action tokens), after which it generates 10 seconds of future gameplay conditioned only on a sequence of held-out controller actions. The generated rollout is then split into 14-frame chunks. This setup enables a comprehensive analysis of the universe’s evaluation capabilities across two axes: (i) in-domain performance: evaluating on Skygarden (Environment A), the environment used for fine-tuning; (ii) generalization: assessing performance on six unseen environments (Environments B–G). It also allows comparison across generation quality and model capacity. We generate 82 rollouts for each model-environment setting, resulting in 656 rollouts in total.
@@
Rollout Filtering. To ensure quality and clarity, we filtered out rollouts that: (i) had no visible agents, (ii) featured stationary agents, (iii) were taken from early uninformative environment segments, or (iv) had significant visual obstruction. We also excluded sequences containing more than four characters to reduce annotation ambiguity.
@@
universe Response Generation. To obtain responses from universe, we provide it with a video segment (resized to match the evaluator’s input resolution) along with its corresponding question. We then sample five responses using greedy decoding and select the most frequent response as the final answer. In cases where all five responses are unique (i.e., no majority), one response is selected at random. The resulting dataset comprises rollouts from 8 model–environment pairs: rollouts generated by WHAM 140M on Environment A (Skygarden), and rollouts generated by WHAM 1.6B across seven distinct environments (Environments A–G). For each model–environment pair, we sample 30 rollouts. Each rollout is annotated with 6 question–answer (QA) pairs, along with a corresponding response from the adapted evaluator. Each of the resulting 1,440 QA instances was rated by 3 annotators, yielding 4,320 total human judgments.
E.2 Evaluation Metrics
@@
Model Env. Valid QA Pairs Cohen’s $\kappa$
WHAM 140M A 24 0.79
WHAM 1.6B A 29 0.91
B 28 0.67
C 28 0.74
D 29 0.87
E 30 0.67
F 29 0.61
G 30 0.59
Sample Size Justification. We annotate 30 rollouts per model–environment pair. Assuming a standard deviation of $\sigma\approx 0.2$ and a 95% confidence level, the confidence interval (CI) width is given by CI $\text{CI Width}=z_{\frac{1-C}{2}}\cdot\frac{\sigma}{\sqrt{n}}$. This yields an estimated CI of $\sim$7.1for individual model–environment pairs ($n=30$), and $\sim$2.5when aggregating across all eight pairs ($n=240$), offering sufficient precision for comparative evaluation.
@@
Model Env. Binary MC OE Binary MC OE
WHAM 140M A 92.9 / 89.3 35.7 / 32.1 41.1 / 39.3 85.7 / 85.7 10.7 / 10.7 60.7 / 60.7
WHAM 1.6B A 98.3 / 96.7 51.7 / 46.7 75.0 / 73.3 93.3 / 93.3 83.3 / 83.3 93.3 / 93.3
B 96.7 / 96.7 60.0 / 60.0 65.0 / 60.0 99.9 / 99.9 90.0 / 90.0 93.3 / 93.3
C 96.7 / 96.7 63.3 / 63.3 80.0 / 80.0 99.9 / 99.9 86.7 / 86.7 93.3 / 93.3
D 93.3 / 93.3 43.3 / 43.3 73.3 / 73.3 96.7 / 96.7 96.7 / 96.7 99.9 / 99.9
E 80.0 / 76.7 76.7 / 73.3 93.3 / 93.3 96.7 / 96.7 99.9 / 99.9 99.8 / 99.8
F 71.7 / 70.0 56.7 / 56.7 75.0 / 70.0 96.7 / 96.7 93.3 / 93.3 96.7 / 96.7
G 68.3 / 66.7 50.0 / 46.7 80.0 / 76.7 93.3 / 93.3 90.0 / 90.0 96.7 / 96.7
E.3 Results
[…]
@@ E.3 Results
Figure 11: Representative frames from rollouts generated across seven environments. WHAM 140M (top row) was trained only on Skygarden (Environment A); WHAM 1.6B (rows 2-8) generalizes across seven environments.
Appendix F Supplementary Experimental Results
@@
F.1 Zero-Shot Performance of PaliGemma Models
@@ F.2 CLIPScore Comparisons
F.3 Low-Rank Adaptation Comparisons

### v2

@@ E.1.3 Prompt Format
[…]
@@
E.1.4 Pretraining Data and Filtering
PaliGemma is pretrained on a mixture of large-scale vision-language datasets, including WebLI (Chen et al., 2022b), CC3M-35L (Sharma et al., 2018), VQ2A-CC3M-35L (Changpinyo et al., 2022), OpenImages (Piergiovanni et al., 2022), and WIT (Srinivasan et al., 2021). Data quality and safety are maintained through pornographic content filtering, text safety and toxicity filtering, and privacy-preserving measures.
E.2 Evaluation Metrics
@@
Formally, let $G$ and $GT$ denote the sets of $n$-grams in the generated and ground truth answers, respectively. Precision and recall are defined as:
$P=\frac{|G\cap GT|}{|G|},\quad R=\frac{|G\cap GT|}{|GT|}$ (3)
where $|G\cap GT|$ counts overlapping $n$-grams. The ROUGE score is then computed as:
@@ E.3 Results
Optimizer AdamW Loshchilov & Hutter (2019)
@@
VLM backbone PaliGemma 2 (3B) Beyer et al. (2024)
Hyperparameters. Table 3 summarizes the core training hyperparameters used across all adaptation experiments. We train all models on 8 NVIDIA A100 GPUs with a batch size of 1 per device and accumulate gradients over 4 steps, yielding an effective batch size of 32. Each epoch corresponds to a full pass over the adaptation dataset, and no early stopping is applied. Models were trained for 1–10 epochs depending on task and setting. Optimization is performed using AdamW (Loshchilov & Hutter, 2019) with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, a base learning rate of $5\times 10^{-5}$, and weight decay of $1\times 10^{-6}$. We use cosine learning rate annealing (Loshchilov & Hutter, 2022) with a linear warmup over the first 10% of training steps. To stabilize training, we apply gradient clipping with a global norm threshold of 1.0. All models use PaliGemma 2 (3B) (Beyer et al., 2024) as the vision-language backbone unless otherwise noted. We vary the number of input frames between 1 and 8 depending on task, and all images are resized to a fixed resolution of $224\times 224$. Training is conducted in bfloat16 precision using data parallelism. Model selection is based on final validation accuracy.
@@
Appendix F Human Annotation Study
@@ F.1 Study Design
Selected World Models. For our study, we aim to evaluate rollouts generated across different model scales, training diversities, and output resolutions, while keeping the underlying architecture general enough to apply broadly. To this end, we select two autoregressive world models (Kanervisto et al., 2025). The autoregressive formulation offers a flexible and widely adopted framework, and is the basis for many state-of-the-art private world models.
The selected models generate sequences of visual frames and controller actions without textual supervision, using a decoder-only transformer (Radford et al., 2019; Vaswani et al., 2017) trained autoregressively on discrete tokens. Visual frames are encoded with a VQGAN (Esser et al., 2021), while joystick actions are tokenized using a learned discretization scheme based on action bucketization (Kanervisto et al., 2020).
@@
Each world model is implemented as a decoder-only transformer (Radford et al., 2019; Vaswani et al., 2017), trained to predict discrete tokens representing visual observations and actions. Visual frames are first compressed with a VQGAN (Esser et al., 2021), while joystick actions are tokenized using a learned discretization scheme based on action bucketization (Kanervisto et al., 2020). The objective is next-token prediction conditioned on prior visual and action tokens. We focus on the following model variants that differ in capacity, training diversity, and output resolution: (i) a large-scale model: 1.6B parameters, trained for 200K steps on gameplay from seven diverse environments (including Skygarden) at $300\times 180$, and (ii) a smaller model: 140M parameters, trained for 100K steps on gameplay from a single environment (Skygarden) at $128\times 128$ resolution.
Rollouts Generation. Rollout generation follows a consistent protocol for both world models: at inference time, the model is conditioned on 1 second of ground-truth gameplay (visual and action tokens), after which it generates 10 seconds of future gameplay conditioned only on a sequence of held-out controller actions. For rollout generation, we use reference trajectories from the game dataset, which include time-aligned controller inputs and image frames. For each generated rollout, we randomly sample a 1-second context window (video frames and corresponding action tokens) from these trajectories. The world model is then conditioned on this context, which consists of an interleaved sequence of tokenized images and actions, and proceeds autoregressively to generate latent image tokens and discretized action tokens. The generated rollout is then split into 14-frame chunks. This setup enables a comprehensive analysis of the universe’s evaluation capabilities across two axes: (i) in-domain performance: evaluating on Skygarden, the environment used for fine-tuning; (ii) generalization: assessing performance on six unseen environments. It also allows comparison across generation quality and model capacity. We generate 82 rollouts for each model-environment setting, resulting in 656 rollouts in total.
Figure 12: Randomly sampled examples of rollouts excluded from the human evaluation study.
Rollout Filtering. To keep evaluation focused on challenging sequences, we remove only those rollouts that are either uninterpretable or uninformative. Specifically, annotators discard: (i) early segments with no meaningful interactions; (ii) rollouts dominated by occlusion; and (iii) sequences with no visible agents. These criteria target cases where neither humans nor models can extract actionable dynamics, while retaining all visually imperfect, ambiguous, and failure-prone generations. Thus, filtering eliminates only sequences that provide no evaluable signal. Figure 12 shows randomly sampled examples from each excluded category.
Figure 13: Randomly sampled examples of rollouts retained for human evaluation.
Complexity of Retained Rollouts. Despite filtering out uninterpretable/uninformative sequences, many retained rollouts exhibit visual imperfections that make evaluation challenging. Annotators highlighted recurring generation failures including: (i) skin and appearance drift, where model-generated agent skins diverge from canonical silhouettes; (ii) missing or corrupted assets, such as hoverboards implied only through motion cues; and (iii) HUD-induced occlusion, where oversized or misplaced UI elements block important scene content. These imperfections arise from the underlying world models and introduce ambiguity while still permitting meaningful assessment. Figure 13 illustrates representative failure-prone cases, emphasizing that universe is evaluated on realistic, imperfect rollouts rather than sanitized inputs.
@@
universe’s Response Generation. To obtain responses from universe, we provide it with a video segment (resized to match the evaluator’s input resolution) along with its corresponding question. We then sample five responses using greedy decoding and select the most frequent response as the final answer. In cases where all five responses are unique (i.e., no majority), one response is selected at random. The resulting dataset comprises rollouts from 8 settings: rollouts generated by a smaller model on Skygarden, and rollouts generated by a larger model across seven distinct environments. For each model–environment pair, we sample 30 rollouts. Each rollout is annotated with 6 question–answer (QA) pairs, along with a corresponding response from the adapted evaluator. Each of the resulting 1,440 QA instances was rated by 3 annotators, yielding 4,320 total human judgments.
F.2 Evaluation Metrics
@@
Setting Valid QA Pairs Cohen’s $\kappa$
1 29 0.91
2 28 0.67
3 28 0.74
4 29 0.87
5 30 0.67
6 29 0.61
7 30 0.59
8 24 0.79
Sample Size Justification. We annotate 30 rollouts per model–environment pair. Assuming a standard deviation of $\sigma\approx 0.2$ and a 95% confidence level, the confidence interval (CI) width is given by $\text{CI Width}=z_{\frac{1-C}{2}}\cdot\frac{\sigma}{\sqrt{n}}$. This yields an estimated CI of $\sim$7.1% for individual model–environment pairs ($n=30$), and $\sim$2.5% when aggregating across all eight pairs ($n=240$), offering sufficient precision for comparative evaluation.
@@
Setting Binary MC OE Binary MC OE
1 98.3 / 96.7 51.7 / 46.7 75.0 / 73.3 93.3 / 93.3 83.3 / 83.3 93.3 / 93.3
2 96.7 / 96.7 60.0 / 60.0 65.0 / 60.0 99.9 / 99.9 90.0 / 90.0 93.3 / 93.3
3 96.7 / 96.7 63.3 / 63.3 80.0 / 80.0 99.9 / 99.9 86.7 / 86.7 93.3 / 93.3
4 93.3 / 93.3 43.3 / 43.3 73.3 / 73.3 96.7 / 96.7 96.7 / 96.7 99.9 / 99.9
5 80.0 / 76.7 76.7 / 73.3 93.3 / 93.3 96.7 / 96.7 99.9 / 99.9 99.8 / 99.8
6 71.7 / 70.0 56.7 / 56.7 75.0 / 70.0 96.7 / 96.7 93.3 / 93.3 96.7 / 96.7
7 68.3 / 66.7 50.0 / 46.7 80.0 / 76.7 93.3 / 93.3 90.0 / 90.0 96.7 / 96.7
8 92.9 / 89.3 35.7 / 32.1 41.1 / 39.3 85.7 / 85.7 10.7 / 10.7 60.7 / 60.7
F.3 Results
Table 15 reports graded and strict accuracy across environments, recognition targets (Action and Character Recognition), and question formats (Binary, Multiple-Choice, Open-Ended). We observe a clear gap in performance between rollouts generated by the two world models. universe struggles with outputs from WHAM 140M, achieving substantially lower accuracy compared to WHAM 1.6B. This is likely due to a mismatch in image resolution: WHAM 140M generates frames at $128\times 128$ resolution, which must be upsampled to the universe’s expected input of $224\times 224$. Despite resizing, the resulting frames often lack sharpness, making actions and characters harder to recognize. In contrast, universe performs well on rollouts from WHAM 1.6B, even across diverse environments. On the in-domain setting (Environment A), the model achieves strong results—averaging 75.02% graded accuracy for AR and 90.00% for CR. When evaluating on the six unseen environments (Environments B–G), performance for AR drops slightly (from 75.02% to 73.52%), while CR remains stable or improves, suggesting strong generalization in character grounding and visual consistency tracking.
@@ F.3 Results
Figure 14: Representative frames from rollouts across the eight evaluation settings, spanning different environments, scales, and resolutions.
Appendix G Supplementary Experimental Results
@@
G.1 GPT-5 Performance
Figure 15: Performance of GPT-5 on six randomly sampled binary action recognition questions. Each panel shows the trial prompt, the model’s response, and the ground-truth label. Despite the apparent simplicity of the setup, GPT-5 fails on 5/6 trials, underscoring the difficulty of the task and the need for task-specific adaptation.
To demonstrate the complexity of the tasks that comprise our protocol, we conducted an evaluation of GPT-5 on randomly selected examples. We deliberately chose the simplest evaluation regime (binary action recognition) to test whether the model can succeed without adaptation.
Figure 15 demonstrates that out of six random samples, GPT-5 produced incorrect answers in five cases. This consistent failure highlights the difficulty of the task.
G.2 Zero-Shot Performance of PaliGemma Models
@@ G.3 CLIPScore Comparisons
G.4 Low-Rank Adaptation Comparisons
@@
Appendix H Comparison with Existing Evaluation Baselines
To validate that universe captures evaluation dimensions beyond existing automated metrics, we analyze its correlation with two representative baselines: Fréchet Video Distance (FVD), a standard distributional metric for video generation quality, and VBench, a comprehensive multi-dimensional benchmark for text-to-video evaluation. These comparisons assess whether universe’s focus on action-conditioned semantic alignment provides complementary signal to metrics designed primarily for perceptual quality and temporal coherence.
H.1 Comparison with FVD
Table 18: Pearson correlation between FVD and universe across experimental settings. None of the correlations are statistically significant (all $p>0.05$).
Setting Action Recognition Character Recognition
1 0.09 0.03
2 -0.07 -0.08
3 -0.17 0.06
4 -0.03 -0.10
5 0.07 -0.25
6 -0.01 0.21
7 0.32 0.03
8 -0.10 0.33
FVD Unterthiner et al. (2018) is a widely used reference-free metric for evaluating generative video models, defined as a distance between feature distributions of real and generated videos. Although commonly treated as a default quantitative proxy for rollout quality in world-model evaluation, FVD is agnostic to task semantics and action alignment—the dimensions universe is specifically designed to capture. Across the eight environments, raw FVD values are tightly clustered (means between 3.23 and 3.27 with standard deviations below 0.05), reflecting limited dynamic range and strong dataset-dependent scale effects. As a result, raw scores are not directly comparable across settings, making correlations a more informative basis for analysis. We therefore assess the relationship between the two metrics by computing Pearson correlations between FVD and universe’s human-aligned scores across all environments for both Action Recognition and Character Recognition tasks.
Results. For each environment (Settings 1–8) and each task, we obtain FVD scores on the same set of rollouts evaluated by universe, report the scores and compute the correlation between the two. The resulting coefficients are reported in Table 18. Correlations are small in magnitude and unstable in sign, ranging from 0.25-0.25 to $0.33$, with most values close to zero. The largest positive correlations are modest (Action Recognition, Setting 7: $r=0.32$; Character Recognition, Setting 8: $r=0.33$), and several settings exhibit weak negative correlations (e.g., Character Recognition, Setting 5: r=0.25r=-0.25). None of these correlations are statistically significant ($p>0.05$ for all settings), indicating universe captures complementary semantic cues. These results demonstrate that FVD and universe provide complementary evaluation signals: FVD captures distribution-level perceptual quality and temporal coherence, while universe quantifies task-specific semantic alignment through human-grounded assessment.
H.2 Correlation with VBench Metrics
Figure 17: Correlation between universe and VBench evaluation suite across eight environments and six dimensions. The human action dimension is omitted as it was constant and non-informative. Correlations remain consistently low to moderate (|r| < 0.4), with none of the correlations are statistically significant (all $p>0.05$).
To validate that universe captures evaluative dimensions beyond existing automated benchmarks, we compare its human-aligned evaluation scores against VBench Huang et al. (2024), a comprehensive state-of-the-art benchmark for generative video evaluation. We computed correlations on identical rollout sets across eight experimental environments using six evaluation configurations, parameterized by task (Action Recognition or Character Recognition) and prompt type (binary, multiple-choice, open-ended). The analysis spans seven VBench dimensions: Temporal Flickering, Temporal Style, Background Consistency, Overall Consistency, Subject Consistency, Dynamic Degree, and Human Action. The Human Action dimension is omitted from visualization as it exhibited no variance across samples.
Results. Figure 17 shows that correlations between universe and VBench metrics remain consistently low to moderate ($|r|<0.4$) across all valid comparisons. Average correlations range from -0.037 for Subject Consistency to 0.06 for Temporal Flickering, with most values falling within 0.1-0.4. Notably, the sign variability across settings (including both positive and negative correlations) indicates that universe captures complementary semantic cues not reflected in VBench’s quality dimensions. This divergence is expected: while VBench focuses on low-level perceptual quality and temporal coherence, universe explicitly targets semantic alignment.
Summary. The consistent orthogonality observed across both analyses demonstrates that universe provides a complementary evaluation dimension to existing benchmarks. While standard distributional metrics (FVD) capture visual fidelity and text-to-video benchmarks (VBench) assess perceptual quality, universe quantifies action-conditioned semantic alignment through human-grounded evaluation. Comprehensive world model assessment therefore benefits from combining these approaches: existing metrics for perceptual coherence and universe for task-relevant semantic fidelity.